OpenStudy (anonymous):

How many solutions can be found for the equation 4x − 8 − 4x = 9? Zero One Two Infinitely many

OpenStudy (anonymous):

is it indefinitely many

OpenStudy (anonymous):

why would you think that?

OpenStudy (anonymous):

because it has 2 4xs?

OpenStudy (anonymous):

what did you get as the simplified answer?

OpenStudy (anonymous):

idk how to solve it

OpenStudy (anonymous):

and how does that make it infinite?

OpenStudy (anonymous):

wow. ok.

OpenStudy (anonymous):

im just guessing

OpenStudy (anonymous):

so you +8 on both sides forst

OpenStudy (anonymous):

17

OpenStudy (anonymous):

then simplify the x'es

OpenStudy (anonymous):

idk how to do that part

OpenStudy (anonymous):

well... what do you get as the equation after u +8?

OpenStudy (anonymous):

4x - 4x = 17? idk im very confused

OpenStudy (anonymous):

well that's actually right...

OpenStudy (anonymous):

now what's 4x-4x?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

right... so what is the equation now?

OpenStudy (anonymous):

0= 17

OpenStudy (anonymous):

so the answer is 0?

OpenStudy (anonymous):

does that have a solution?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

is there a x in the equation...?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

ok. so what is the answer for the problem then?

OpenStudy (anonymous):

you are my savior thanks so much. and zero..

OpenStudy (anonymous):

lol :P np

OpenStudy (anonymous):

btw what grade level math is this?

OpenStudy (anonymous):

8th pre algebra

OpenStudy (anonymous):

I GOT 100 THANK YOU

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

good for you

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 4x- 8- 4x = 9 }\) regardless of the value of x, we would simplify the equation to subtract the 4x's \(\large\color{black}{ \displaystyle \cancel{4x}- 8\cancel{- 4x} = 9 }\) \(\large\color{black}{ \displaystyle - 8= 9 }\) And this is always false (again - no matter what we set x equal to, because as there is no x in the equation we are not plugging anything in, and thus, for every value of x, we will be getting that -8=9 -- and that is false)